# Graph Alignment via Dual-Pass Spectral Encoding and Latent Space Communication

### v1

@@ Appendix B Proof of Theorem 1
Preliminary Definitions: Let GCN filter $\tilde{\mathbf{A}}_{\text{GCN},\text{sym}}=\mathbf{I}-\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}(\mathbf{I}-\tilde{\mathbf{\Lambda}})\mathbf{U}^{\top}$, where $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{n}]$ is the eigenbasis, $\boldsymbol{\hat{\Lambda}}=\text{diag}(\hat{\lambda}_{1},\ldots,\hat{\lambda}_{n})$ is the diagonal matrix of eigenvalues, and each eigenvalue satisfies $\hat{\lambda}_{i}\in[0,2]$. The spectral representation of node features is: $\mathbf{X}=\sum_{k=1}^{n}\hat{\mathbf{X}}_{k}\mathbf{u}_{k}$, where $\hat{\mathbf{X}}_{k}=\mathbf{u}_{k}^{\top}\mathbf{X}$.
  • The low-pass component captures the smooth, global structure of the graph. It aggregates information from neighbors, producing embeddings: $\mathbf{Z}_{\text{low}}=\sum_{k=1}^{n}p_{\text{low}}(\hat{\lambda}_{k})\,\hat{\mathbf{X}}_{k}\,\mathbf{u}_{k}$, where $p_{\text{low}}(\hat{\lambda}_{k})=1-\frac{1}{2}\hat{\lambda}_{k}$. This captures smoothed signals over the graph—node embeddings are averages of their neighbors.
  • The high-pass component captures complementary, local variations and finer structural details, given by: $\mathbf{Z}_{\text{high}}=\sum_{k=1}^{n}p_{\text{high}}(\hat{\lambda}_{k})\,\hat{\mathbf{X}}_{k}\,\mathbf{u}_{k}$, where $p_{\text{high}}(\hat{\lambda}_{k})=\frac{1}{2}\hat{\lambda}_{k}$.
Claim 1 (Neighborhood preservation). The dual-pass embedding $\mathbf{z}_{i}$ preserves neighborhood similarity as effectively as the low-pass embedding $\mathbf{z}_{i}^{\text{low}}$.
@@
$\langle\mathbf{z}_{i},\mathbf{z}_{j}\rangle=\langle\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{low}}\rangle+\langle\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{high}}\rangle+\langle\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{high}}\rangle+\langle\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{low}}\rangle.$
@@
$p_{\text{low}}(\hat{\lambda}_{k})+p_{\text{high}}(\hat{\lambda}_{k})=(1-\frac{1}{2}\hat{\lambda}_{k})+\frac{1}{2}\hat{\lambda}_{k}=1,\quad\forall\hat{\lambda}_{k}\in[0,2].$
This spectral separation implies that $\mathbb{E}\big{[}\langle\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{i}^{\text{high}}\rangle\big{]}\approx 0$, meaning that the low-pass and high-pass components are approximately orthogonal in expectation.
Therefore, the low-pass and high-pass components are approximately orthogonal, and the high-pass component adds complementary information without significantly affecting the neighborhood-preserving properties encoded by the low-pass component. Consequently, the dual-pass embedding preserves neighborhood similarity as effectively as the low-pass component alone.
Claim 2 (Enhanced discriminability for node correspondence). For node correspondence tasks, the dual-pass embedding $\mathbf{z}_{i}$ provides superior discriminability compared to either $\mathbf{z}_{i}^{\text{low}}$ or $\mathbf{z}_{i}^{\text{high}}$ alone. That is, false correspondences are less likely under similarity computed via $\mathbf{z}_{i}$.
@@
As we mentioned, the dual-pass filter design ensures perfect spectral complementarity: the low-pass and high-pass filters have anti-correlated frequency responses, ensuring that their contributions are nearly independent. Consequently, their mutual information is bounded, $I(\mathbf{z}_{i}^{\text{low}};\mathbf{z}_{i}^{\text{high}})\leq\epsilon$, for small $\epsilon>0$, reflecting the opposing spectral emphasis.
@@
$d(\mathbf{z}_{i},\mathbf{z}_{j})^{2}=d(\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{low}})^{2}+d(\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{high}})^{2}+2\langle\mathbf{z}_{i}^{\text{low}}-\mathbf{z}_{j}^{\text{low}},\mathbf{z}_{i}^{\text{high}}-\mathbf{z}_{j}^{\text{high}}\rangle.$
@@
$d(\mathbf{z}_{i},\mathbf{z}_{j})^{2}\approx d(\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{low}})^{2}+d(\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{high}})^{2}.$
The separation margin is thus defined as $\Delta=\min_{(i,j)\in\mathcal{C}}\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})-\max_{(i,k)\in\mathcal{N}}\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})$.
@@
  • Case A (Low-pass insufficient): When graphs share similar global structure but differ in local details, $\mathbf{z}_{i}^{\text{low}}$ may yield high similarity for non-corresponding pairs. However, $\mathbf{z}_{i}^{\text{high}}$ captures local differences, reducing false positives.
  • Case B (High-pass insufficient): When local structures are noisy or similar, $\mathbf{z}_{i}^{\text{high}}$ may be unreliable. However, $\mathbf{z}_{i}^{\text{low}}$ provides stable global discrimination based on community structure and smooth attributes.
@@
$\Delta_{\text{dual}}\geq\max(\Delta_{\text{low}},\Delta_{\text{high}})+\gamma,$
@@ C.1 Benchmarks
  • Celegans: This dataset models the protein-protein interaction network of Caenorhabditis elegans. Each node represents a protein, and edges indicate physical or functional interactions between proteins, making it useful for biological network analysis and alignment tasks involving molecular networks [Kunegis, 2013].
  • Arenas: A communication network derived from email exchanges at the University Rovira i Virgili. Nodes correspond to individual users, and edges represent the presence of at least one email sent between them. It serves as a social interaction graph with temporal and communication patterns [Leskovec and Krevl, 2014].
  • Douban: A social network from the Chinese movie review platform Douban, where nodes represent users, and edges capture friend or contact relationships. This dataset is commonly used to study social dynamics and network alignment in social media contexts [Zhang and Tong, 2016].
  • Cora: A citation network of scientific papers where nodes are publications, and edges denote citation relationships. Cora is a benchmark dataset for graph mining and node classification, providing a structured academic citation graph ideal for evaluating graph-based learning models [Sen et al., 2008].
  • DBLP: An extensive citation network aggregated from DBLP, Association for Computing Machinery (ACM), Microsoft Academic Graph (MAG), and other scholarly databases. It includes publication and citation information, widely used for testing graph alignment, clustering, and knowledge discovery tasks in academic networks [Pan et al., 2016].
  • CoauthorCs: A co-authorship network in computer science that represents collaborations between authors. Nodes correspond to researchers, and edges indicate joint publications. It is often used to study community structure and author disambiguation in bibliographic databases [Sinha et al., 2015].
  • ACM-DBLP: This dataset contains two co-authorship graphs from the ACM and DBLP databases. Nodes represent authors, and edges indicate co-authorship. Although collected independently, both graphs share overlapping authors, with 6,325 ground-truth alignments. Node features capture publication distributions across research venues. The dataset poses a challenging alignment task due to structural and feature discrepancies between the two graphs [Zhang and Tong, 2018].
  • Douban Online-Offline: This dataset comprises two social graphs from the Douban platform—one based on online interactions and the other on offline event co-attendance. Both graphs share a subset of users, with 1,118 aligned nodes. Node features reflect user location distributions. The dataset is designed to evaluate alignment across heterogeneous and partially overlapping social networks [Zhang and Tong, 2016].
@@ C.2 Performance metrics
To assess graph alignment performance, we adopt two widely used metrics: alignment accuracy (Acc) and Hit@$k$. Alignment Accuracy (Acc) measures the proportion of correctly predicted node correspondences among all ground-truth aligned pairs, providing a direct measure of overall matching performance. Hit@$k$ evaluates whether the true corresponding node from the source graph appears within the top-$k$ predicted candidates for each node in the target graph. This metric reflects the ability of model to rank correct matches highly and is particularly useful for top-$k$ retrieval scenarios.
Both metrics are computed using all available ground-truth node pairs, with higher values indicating better alignment quality.
C.3 Implementation details
As shown in Figure 2, our proposed architecture consists of two GAEs, each equipped with a dual-pass GCN encoder. The number of GCN layers is selected based on empirical performance for each benchmark: 2 layers for ACM-DBLP and the semi-synthetic datasets, 3 for DBLP, 5 for Douban, and 6 for Douban Online-Offline.
@@ C.4.1 Setup
We evaluate the vision-language alignment task on a range of benchmarks, including CIFAR-10 [Krizhevsky et al., 2009], CINIC-10 [Darlow et al., 2018], CIFAR-100 [Krizhevsky et al., 2009], and ImageNet-100 [Russakovsky et al., 2015], using representations extracted from diverse pretrained vision and language models. For each vision model, class-level representations are derived by averaging image-level embeddings within each class. Correspondingly, language representations are obtained by averaging embeddings generated from multiple textual prompts for each class. To enable application of our graph alignment method, we build a similarity graph from these representations, where each class-level embedding is treated as a node and connected to its $k$ most similar neighbors according to cosine similarity.
C.4.2 Baselines
@@
We compare against a set of established solvers and heuristics for the alignment problem. LocalCKA [Maniparambil et al., 2024] leverages the centered kernel alignment (CKA) metric to approximate the QAP with a linear assignment formulation, providing an efficient method for vision–language correspondence. Optimal Transport (OT) methods [Peyré et al., 2016] address the alignment by modeling embeddings as probability distributions and computing the minimal transport cost, thereby preserving geometric structure across modalities. The Fast Approximate QAP algorithm (FAQ) [Vogelstein et al., 2015] is a well-known primal heuristic that relaxes the QAP and iteratively refines the solution, yielding scalable but approximate alignments. MPOpt [Hutschenreiter et al., 2021] represents a generic mathematical programming approach, solving the alignment as a constrained optimization problem using standard formulations. Gurobi [Gurobi Optimization, LLC, 2023] is a commercial off-the-shelf solver for mixed-integer and quadratic programs, providing near-optimal results for small problem instances. Finally, the Hahn-Grant solver [Schnaus et al., 2025] is a dual ascent algorithm that produces strong lower bounds by repeatedly solving linear assignment problems.
C.4.3 Vision and language models

### v2

@@ Appendix B Proof of Theorem 1
Preliminary Definitions: Let GCN filter $\tilde{\mathbf{A}}_{\mathrm{GCN},\mathrm{sym}}=\mathbf{I}-\tilde{\mathbf{L}}_{\mathrm{sym}}=\mathbf{U}(\mathbf{I}-\tilde{\mathbf{\Lambda}})\mathbf{U}^{\top}$, where $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{n}]$ is the eigenbasis, $\bm{\hat{\Lambda}}=\mathrm{diag}(\hat{\lambda}_{1},\ldots,\hat{\lambda}_{n})$ is the diagonal matrix of eigenvalues, and each eigenvalue satisfies $\hat{\lambda}_{i}\in[0,2]$. The spectral representation of node features is: $\mathbf{X}=\sum_{k=1}^{n}\hat{\mathbf{X}}_{k}\mathbf{u}_{k}$, where $\hat{\mathbf{X}}_{k}=\mathbf{u}_{k}^{\top}\mathbf{X}$.
  • The low-pass component captures the smooth, global structure of the graph. It aggregates information from neighbors, producing embeddings: $\mathbf{Z}_{\mathrm{low}}=\sum_{k=1}^{n}p_{\mathrm{low}}(\hat{\lambda}_{k})\,\hat{\mathbf{X}}_{k}\,\mathbf{u}_{k}$, where $p_{\mathrm{low}}(\hat{\lambda}_{k})=1-\frac{1}{2}\hat{\lambda}_{k}$. This captures smoothed signals over the graph, node embeddings are averages of their neighbors.
  • The high-pass component captures complementary, local variations and finer structural details, given by: $\mathbf{Z}_{\mathrm{high}}=\sum_{k=1}^{n}p_{\mathrm{high}}(\hat{\lambda}_{k})\,\hat{\mathbf{X}}_{k}\,\mathbf{u}_{k}$, where $p_{\mathrm{high}}(\hat{\lambda}_{k})=\frac{1}{2}\hat{\lambda}_{k}$.
Claim 1 (Neighborhood preservation). The dual-pass embedding $\mathbf{z}_{i}$ preserves neighborhood similarity as effectively as the low-pass embedding $\mathbf{z}_{i}^{\mathrm{low}}$.
@@
$\langle\mathbf{z}_{i},\mathbf{z}_{j}\rangle=\langle\mathbf{z}_{i}^{\mathrm{low}},\mathbf{z}_{j}^{\mathrm{low}}\rangle+\langle\mathbf{z}_{i}^{\mathrm{high}},\mathbf{z}_{j}^{\mathrm{high}}\rangle+\langle\mathbf{z}_{i}^{\mathrm{low}},\mathbf{z}_{j}^{\mathrm{high}}\rangle+\langle\mathbf{z}_{i}^{\mathrm{high}},\mathbf{z}_{j}^{\mathrm{low}}\rangle.$
@@
$p_{\mathrm{low}}(\hat{\lambda}_{k})+p_{\mathrm{high}}(\hat{\lambda}_{k})=(1-\frac{1}{2}\hat{\lambda}_{k})+\frac{1}{2}\hat{\lambda}_{k}=1,\quad\forall\hat{\lambda}_{k}\in[0,2].$
Due to their complementary spectral responses, the low-pass and high-pass components are approximately orthogonal. To see this, note that the low-pass filter $p_{\text{low}}(\hat{\lambda}_{k})=1-\frac{1}{2}\hat{\lambda}_{k}$ is monotonically decreasing, achieving maximum response at $\hat{\lambda}_{k}=0$ and minimum at $\hat{\lambda}_{k}=2$. Conversely, the high-pass filter $p_{\text{high}}(\hat{\lambda}_{k})=\frac{1}{2}\hat{\lambda}_{k}$ is monotonically increasing, with minimum response at $\hat{\lambda}_{k}=0$ and maximum at $\hat{\lambda}_{k}=2$. The spectral overlap between components is measured by the product $p_{\text{low}}(\hat{\lambda}_{k})\cdot p_{\text{high}}(\hat{\lambda}_{k})=\frac{1}{4}\hat{\lambda}_{k}(2-\hat{\lambda}_{k})$, which is maximized only at the intermediate eigenvalue $\hat{\lambda}_{k}=1$ and approaches zero at both extremes.
This spectral disjointness ensures that the high-pass component adds complementary discriminative information without interfering with the neighborhood-preserving properties encoded in the low-frequency domain by the low-pass component. Consequently, the dual-pass embedding $\mathbf{z}_{i}=[\mathbf{z}_{i}^{\text{low}}\,\|\,\mathbf{z}_{i}^{\text{high}}]$ preserves neighborhood similarity as effectively as the low-pass component $\mathbf{z}_{i}^{\text{low}}$ alone, since the neighborhood-relevant information is fully retained while additional discriminative power is gained.
Claim 2 (Enhanced discriminability for node correspondence). For node correspondence tasks, the dual-pass embedding $\mathbf{z}_{i}$ provides superior discriminability compared to either $\mathbf{z}_{i}^{\mathrm{low}}$ or $\mathbf{z}_{i}^{\mathrm{high}}$ alone. That is, false correspondences are less likely under similarity computed via $\mathbf{z}_{i}$.
@@
As we mentioned, the dual-pass filter design ensures perfect spectral complementarity: the low-pass and high-pass filters have anti-correlated frequency responses, ensuring that their contributions are nearly independent. Consequently, their mutual information is bounded, $I(\mathbf{z}_{i}^{\mathrm{low}};\mathbf{z}_{i}^{\mathrm{high}})\leq\epsilon$, for small $\epsilon>0$, reflecting the opposing spectral emphasis.
@@
$d(\mathbf{z}_{i},\mathbf{z}_{j})^{2}=d(\mathbf{z}_{i}^{\mathrm{low}},\mathbf{z}_{j}^{\mathrm{low}})^{2}+d(\mathbf{z}_{i}^{\mathrm{high}},\mathbf{z}_{j}^{\mathrm{high}})^{2}+2\langle\mathbf{z}_{i}^{\mathrm{low}}-\mathbf{z}_{j}^{\mathrm{low}},\mathbf{z}_{i}^{\mathrm{high}}-\mathbf{z}_{j}^{\mathrm{high}}\rangle.$
@@
$d(\mathbf{z}_{i},\mathbf{z}_{j})^{2}\approx d(\mathbf{z}_{i}^{\mathrm{low}},\mathbf{z}_{j}^{\mathrm{low}})^{2}+d(\mathbf{z}_{i}^{\mathrm{high}},\mathbf{z}_{j}^{\mathrm{high}})^{2}.$
The separation margin is thus defined as $\Delta=\min_{(i,j)\in\mathcal{C}}\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{j})-\max_{(i,k)\in\mathcal{N}}\mathrm{sim}(\mathbf{z}_{i},\mathbf{z}_{k})$.
@@
  • Case A (Low-pass insufficient): When graphs share similar global structure but differ in local details, $\mathbf{z}_{i}^{\mathrm{low}}$ may yield high similarity for non-corresponding pairs. However, $\mathbf{z}_{i}^{\mathrm{high}}$ captures local differences, reducing false positives.
  • Case B (High-pass insufficient): When local structures are noisy or similar, $\mathbf{z}_{i}^{\mathrm{high}}$ may be unreliable. However, $\mathbf{z}_{i}^{\mathrm{low}}$ provides stable global discrimination based on community structure and smooth attributes.
@@
$\Delta_{\mathrm{dual}}\geq\max(\Delta_{\mathrm{low}},\Delta_{\mathrm{high}})+\gamma,$
@@ C.1 Benchmarks
  • Celegans: This dataset models the protein-protein interaction network of Caenorhabditis elegans. Each node represents a protein, and edges indicate physical or functional interactions between proteins, making it useful for biological network analysis and alignment tasks involving molecular networks (Kunegis, 2013).
  • Arenas: A communication network derived from email exchanges at the University Rovira i Virgili. Nodes correspond to individual users, and edges represent the presence of at least one email sent between them. It serves as a social interaction graph with temporal and communication patterns (Leskovec & Krevl, 2014).
  • Douban: A social network from the Chinese movie review platform Douban, where nodes represent users, and edges capture friend or contact relationships. This dataset is commonly used to study social dynamics and network alignment in social media contexts (Zhang & Tong, 2016).
  • Cora: A citation network of scientific papers where nodes are publications, and edges denote citation relationships. Cora is a benchmark dataset for graph mining and node classification, providing a structured academic citation graph ideal for evaluating graph-based learning models (Sen et al., 2008).
  • DBLP: An extensive citation network aggregated from DBLP, Association for Computing Machinery (ACM), Microsoft Academic Graph (MAG), and other scholarly databases. It includes publication and citation information, widely used for testing graph alignment, clustering, and knowledge discovery tasks in academic networks (Pan et al., 2016).
  • CoauthorCs: A co-authorship network in computer science that represents collaborations between authors. Nodes correspond to researchers, and edges indicate joint publications. It is often used to study community structure and author disambiguation in bibliographic databases (Sinha et al., 2015).
  • ACM-DBLP: This dataset contains two co-authorship graphs from the ACM and DBLP databases. Nodes represent authors, and edges indicate co-authorship. Although collected independently, both graphs share overlapping authors, with 6,325 ground-truth alignments. Node features capture publication distributions across research venues. The dataset poses a challenging alignment task due to structural and feature discrepancies between the two graphs (Zhang & Tong, 2018).
  • Douban Online-Offline: This dataset comprises two social graphs from the Douban platform, one based on online interactions and the other on offline event co-attendance. Both graphs share a subset of users, with 1,118 aligned nodes. Node features reflect user location distributions. The dataset is designed to evaluate alignment across heterogeneous and partially overlapping social networks (Zhang & Tong, 2016).
@@ C.2 Performance metrics
To assess graph alignment performance, we adopt two widely used metrics: alignment accuracy (Acc) and Hit@$k$. Alignment Accuracy (Acc) measures the proportion of correctly predicted node correspondences among all ground-truth aligned pairs, providing a direct measure of overall matching performance. Hit@$k$ evaluates whether the true corresponding node from the source graph appears within the top-$k$ predicted candidates for each node in the target graph. This metric reflects the ability of model to rank correct matches highly and is particularly useful for top-$k$ retrieval scenarios. Both metrics are computed using all available ground-truth node pairs, with higher values indicating better alignment quality.
C.3 Experimental settings
In section 4.1, we follow the experimental setting introduced in the (He et al., 2024) for generating inconsistent graph pairs. Given a source graph $\mathcal{G}_{s}$ with adjacency matrix $\mathbf{A}$, we construct 10 perturbed and permuted target graphs using the transformation $\hat{\mathbf{A}}=\mathbf{P}(\mathbf{A}+\mathbf{M})\mathbf{P}^{\top}$, where $\mathbf{M}\in\{-1,0,1\}^{N\times N}$ introduces edge-level perturbations, and $\mathbf{P}$ is a random permutation matrix. The perturbation level is controlled by a parameter $p\in\{0,1\%,5\%\}$, representing the fraction of edges modified: $p|\mathcal{E}|$. We adopt seven structural node features from (Berlingerio et al., 2013): node degree, clustering coefficient, average degree of neighbors, average clustering coefficient of neighbors, number of edges in ego-network, number of outgoing edges of ego-network, and number of neighbors of ego-network. These descriptors provide compact, structure-aware representations for robust evaluation under varying structural inconsistencies.
As shown in Figure 2, our proposed architecture consists of two GAEs, each using a dual-pass GCN encoder with shared parameters across both modalities. The number of GCN layers is selected based on empirical performance for each benchmark: 2 layers for ACM-DBLP and the semi-synthetic datasets, 3 for DBLP, 5 for Douban, and 6 for Douban Online-Offline.
@@ C.3 Experimental settings
All other hyperparameters are kept fixed across benchmarks, with values set as follows: $\alpha=10^{-3}$, $\beta=10^{-2}$, $\lambda_{\mathrm{FM}}=1$, $\lambda_{\mathrm{bij}}=10^{-1}$, and $\lambda_{\mathrm{orth}}=10^{-1}$. An ablation study on these hyperparameters is presented in Section D.
@@ C.4.1 Setup
We evaluate the vision-language alignment task on a range of benchmarks, including CIFAR-10 (Krizhevsky et al., 2009), CINIC-10 (Darlow et al., 2018), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet-100 (Russakovsky et al., 2015), using representations extracted from diverse pretrained vision and language models. For each vision model, class-level representations are derived by averaging image-level embeddings within each class. Correspondingly, language representations are obtained by averaging embeddings generated from multiple textual prompts for each class. To enable application of our graph alignment method, we build a similarity graph from these representations, where each class-level embedding is treated as a node and connected to its $k$ most similar neighbors according to cosine similarity.
Since the vision and language models generally produce embeddings of different dimensionalities, in this experiment, we employ dual-pass GCN encoders without weight sharing. While this design accommodates modality-specific feature spaces, it also makes the alignment task more challenging, as the model must learn to reconcile heterogeneous latent representations.
C.5 Hyperparameters
We adopt a similar hyperparameter configuration for the vision-language benchmarks. Specifically, we set $k=5$ when constructing the $k$-NN graphs. Each modality is encoded with a 4-layer dual-pass GCN encoder, and we use 9 Laplacian eigenvectors for CIFAR-10 and CINIC-10, and 90 eigenvectors for CIFAR-100 and ImageNet-100. The hidden and output dimensions of the encoder are both set to 512. All other hyperparameters follow the general settings described in Section C.3.
C.5.1 Baselines
@@
We compare against a set of established solvers and heuristics for the alignment problem. LocalCKA (Maniparambil et al., 2024) leverages the centered kernel alignment (CKA) metric to approximate the QAP with a linear assignment formulation, providing an efficient method for vision–language correspondence. Optimal Transport (OT) methods (Peyré et al., 2016) address the alignment by modeling embeddings as probability distributions and computing the minimal transport cost, thereby preserving geometric structure across modalities. The Fast Approximate QAP algorithm (FAQ) (Vogelstein et al., 2015) is a well-known primal heuristic that relaxes the QAP and iteratively refines the solution, yielding scalable but approximate alignments. MPOpt (Hutschenreiter et al., 2021) represents a generic mathematical programming approach, solving the alignment as a constrained optimization problem using standard formulations. Gurobi (Gurobi Optimization, LLC, 2023) is a commercial off-the-shelf solver for mixed-integer and quadratic programs, providing near-optimal results for small problem instances. Finally, the Hahn-Grant solver (Schnaus et al., 2025) is a dual ascent algorithm that produces strong lower bounds by repeatedly solving linear assignment problems.
We also reference two recent vision–language alignment methods. Vec2Vec (Jha et al., 2025) maps embeddings from different models into a shared latent space using input/output adapters, a shared backbone, and adversarial plus structural losses. CycleReward (Bahng et al., 2025) learns vision–language alignment via cycle-consistency-based preference data and a reward model. These models are not designed for graph alignment and therefore are not direct competitors. Moreover, a full evaluation would require reproducing their results on our benchmarks, which is beyond the scope of this work and is deferred to future studies focused specifically on this domain.
C.5.2 Vision and language models